# MEEG and AT-DGNN: Improving EEG Emotion Recognition with Music Introducing and Graph-based Learning

### v2

@@ VI Conclusion
This study introduces the MEEG dataset and the AT-DGNN framework, marking substantial progress in the field of EEG-based emotion recognition. The MEEG dataset uniquely employs music to induce emotional states, providing an invaluable resource for probing brain responses to emotional stimuli. The AT-DGNN, which integrates attention mechanisms with DGNN, adeptly captures the intricate temporal dynamics of brain activity, thereby markedly enhancing emotion recognition accuracy over existing SOTA methods. This advancement not only propels BCI technology forward but also facilitates novel investigations into the interplay among music, emotion, and cerebral activity. The exemplary performance of AT-DGNN, validated through rigorous comparisons with traditional models, solidifies its utility in advancing neuropsychological understanding.

### v3

@@ VI Conclusion
This study introduces the MEEG dataset and the AT-DGNN framework, significantly advancing EEG-based emotion recognition. The MEEG dataset utilizes music to induce emotional states, providing a unique and critical resource for analyzing brain responses to emotional stimuli. The AT-DGNN framework captures the complex temporal dynamics of brain activity, enhancing emotion recognition accuracy beyond existing SOTA methods. Moreover, the MEEG dataset displays emotional states of subjects with greater precision, enabling SOTA methods to achieve higher accuracy and establishing a more robust benchmark for evaluating EEG-based emotion analysis models. These advancements propel BCI technology forward and facilitate new research into the relationships between music, emotion, and brain activity.
VII ACKNOWLEDGEMENTS
This paper is supported in part by the Shenzhen Fundamental Research Program under Grant JCYJ20230807094104009.